OpenStudy (anonymous):

10. What single discount is equivalent to two successive discounts of 9% and 14%?

OpenStudy (anonymous):

.86(.91x) 78.26% Discount would be 21.74%

OpenStudy (anonymous):

can you please explain how you got this answer. I ended up with something totally different

OpenStudy (dumbcow):

discount of 9% is same as 100% - 9% or 91% of current price assume starting price is 100 after first discount, new price = 100*(1-.09) = 100*.91 = 91 2nd discount of 14% is applied to this new price of 91 final price = 91*(1-.14) = 91*.86 = 78.26 total discount is how much we took away from starting price 100*(1-d) = 78.26 1-d = .7826 d = 1-.7826 d= .2174 or 21.74%